Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

NIght report page - 20131013

Observers: Cosimo Inserra, Ting-Wan Chen
Support team: J. Polshaw, S. Smartt
TAT on duty: Nancy Elias-Rosa, Morgan Fraser

------------------------------------------------

Subrun 3, night 1

Instruments: EFOSC2

WEATHER CONDITION
beginning: wind speed 2.5 m/s; humidity 33 % ; no clouds
middle: wind speed  5.4 m/s; humidity 17 %;  no clouds
end: wind speed  8.2 m/s; humidity 10%;  no clouds 

CALIBRATIONS
EFOSC2 bias
EFOSC2 Spectroscopic flats and rcs
EFOSC2 Skyflats BRz

OBSERVATIONS:
Target                                    Type                 UT         Filter/Grism      Slit            Seeing    Grade (A/B/C)     Comments
LTT7379    STD    23:50    gm11/gm13/gm16    1./1.5    0.9"    A (re-focus)
SN2012ca    FOLLOW-UP    00:03    gm13    1.    0.75"    A
SN2013fs    FOLLOW-UP    (strange noise in the images, so we observed images first, hope it will decrease) B

(UT 02:25 - 02:54) ISSUE TIME
We did a test bias for checking the noise, the pattern was still there and was the same found for LSQ13bnx in the classification spectrum.
The TO at the beginning tried to change a cable at the NTT, but it did not improve, after that we restarted EFOSC2
but even in this case we did not have a good luck. 
After a discussion with the daily TO and that of the 3.6 we decided to change the wind rotator
and apparently it worked. Unfortunately it worked for just 1 image and from the 2nd we had again the problem.
The TO went again to the NTT to physically shut down the detector and restart the machine.
After 4 test clear images without the electronic noise we decided to start again the observation and do not switch to SOFI.

OGLE-2013-SN-079    FOLLOW-UP    03:33    g11/g16    1.    1"    C         (Aborted after the first spectrum)

Changed to SOFI due to the noise pattern (UT 04:35 - 04:44)

OGLE-2013-SN-079    FOLLOW-UP    04:55    JHK    -    0.8"    A
PSNJ01591443+1 90139    FOLLOW-UP    06:12    JHK    -    0.6"    A (We had to do last sky_K after because when we modified the OB, the TO put the wrong number of exposure: 15 instead of 16)
SN2013fc   FOLLOW-UP  08:54  JHK    -    0.7"  A



************************************************************************************************************************************************************************************
Please note that the followup spectra of SN2013fs and OGLE-2013-SN-079 were badly affected by electronic noise and therefore the reductions have not been uploaded to WISeREP.
************************************************************************************************************************************************************************************


Comments